# GrokFormer: Graph Fourier Kolmogorov-Arnold Transformers

### v2

@@ 5 Experiments
In this section, we conduct comprehensive experiments on both synthetic and real-world datasets to verify the effectiveness of our GrokFormer.
@@ 5.1 Performance for Node Classification
Dataset Description. We conduct node classification experiments on 10 widely used datasets in previous graph spectral models (Bo et al., 2023; He et al., 2021), including five homophilic datasets, i.e., Cora, Citeseer, Pubmed, an Amazon co-purchase graph Photo (He et al., 2021), and a co-authorship network Physics (Shchur et al., 2018), and five heterophilic datasets, i.e., Wikipedia graphs Chameleon and Squirrel, the Actor co-occurrence graph (Pei et al., 2020), webpage graphs Texas from WebKB, and Penn94, a large-scale friendship network from the Facebook 100 networks (Lim et al., 2021). A more detailed description can be found in Appendix A.1.
Baselines and Settings. We compare GrokFormer with fifteen competitive baselines, including four spatial-based GNNs, five spectral-based GNNs, and six GTs. Following (He et al., 2021; Bo et al., 2023), we randomly split the node set into train/validation/test set with ratio 60%/20%/20%, and generate 10 random splits to evaluate all models on the same splits. We report the average classification accuracy and standard deviation for each model. For polynomial GNNs, we set the order of polynomials $K$ = 10, consistent with their original setting. For the baseline models, we adopt the best hyperparameters provided by the authors. In the large-scale datasets Physics and Penn94, we implement truncated spectral decomposition for both GrokFormer and Specformer to enhance scalability, selecting the 3,000 eigenvectors associated with the smallest (low-frequency) and largest (high-frequency) eigenvalues. More detailed settings can be found in Appendix A.2.
@@
Results. The results are reported in Table 2, which shows the superiority of GrokFormer over state-of-the-art baselines in both homophilic and heterophilic datasets. Except for the vanilla Transformer, which does not consider graph topology, almost all baselines perform well on homophilic datasets. This is because graph models can easily capture similarity information between neighbors, and low-pass filtering is easy to fit in homophilic networks.
@@ 5.2 Performance for Graph Classification
Dataset. We also conduct graph classification experiments on five datasets from diverse domains. They include three bioinformatics graph datasets, i.e., PROTEINS (Borgwardt et al., 2005), PTC-MR (Toivonen et al., 2003), and MUTAG (Debnath et al., 1991) and two social network datasets, i.e., IMDB-BINARY and IMDB-MULTI (Yanardag & Vishwanathan, 2015) (see Appendix A.1 for more details).
@@ 5.3.1 Pre-defined Filters in Synthetic Datasets
Following prior works (He et al., 2021; Bo et al., 2023; Ma et al., 2024), we generate datasets with six filter patterns of various levels of difficulty. Specifically, images with a resolution of 100 $\times$ 100 from the Image Processing in Matlab library [^1] https://ww2.mathworks.cn/products/image.html are taken, with each image represented as a 2D regular grid graph with 4-neighborhood connectivity. The pixel values serve as node signals ranging from 0 to 1. The $m$-th image has an adjacency matrix $\mathbf{A}_{m}\in\mathbb{R}^{10000\times 10000}$ and a node signal $x_{m}\in\mathbb{R}^{10000}$. For each image, we apply six different predefined filters to the spectral domain of its signal, with each filter detailed in Table 4, where $h_{\delta}(\lambda)$ of Low-comb is defined as $I_{[0,0.5]}(\lambda)+\left|sin(\pi\lambda)\right|I_{(0.5,1)}+\left|sin(2\pi\lambda)\right|I_{[1,2]}$, with $I_{\Omega}=1$ when $\lambda\in\Omega$, $I_{\Omega}=0$ otherwise.
We compare the capability of our GrokFormer filter with six baselines, including GCN, GAT, ChebyNet, GPRGNN, BernNet, and Specformer, in fitting these pre-defined filter patterns through a node regression task. We set the hyper-parameters of the baselines suggested by their authors. For our model, the hyperparameters were probed in $M\in\{32,64,128,256,512\}$, $K\in\{1,2,\cdots,10\}$, and the number of hyperparameters we tuned is not greater than that of the baselines. Two popular evaluation criteria – the sum of squared errors and the $R^{2}$ score – are used.
@@ 5.3.2 Unknown Filters in Real-world Homophilic and Heterophilic Datasets
Adaptability in Graph Spectrum. To analyze the importance of spectrum adaptability, we plot the filters learned by GrokFormer on one typical homophilic dataset, Cora, and one heterophilic dataset, Squirrel, in Figure LABEL:fig:learned_filters. We observe that GrokFormer learns a low-pass filter for Cora, which matches its strong homophily. On Squirrel, which has dense and mostly heterophilic edges, GrokFormer effectively learns the heterophily, resulting in a comb-alike filter. Importantly, GrokFormer does not require prior knowledge to manually tune the spectrum hyperparameters to achieve this; it learns to adaptively fit different filters hidden in the graphs with different homophilic and heterophilic properties. Similar results can be found for the other datasets in Appendix B.2.
Adaptability in Graph Order. Similarly, we also analyze the order adaptability of GrokFormer on Cora and Squirrel. To this end, we evaluate the accuracy performance of GrokFormer with all other settings fixed except that we increase the order $K$ from one to six, from which we observe the relation of $k$ and $K$ w.r.t. the best accuracy.
The results are shown in Figure LABEL:fig:k. Cora is a strong homophilic network that expects a low-pass filter, and such filter is easy to learn, on which GrokFormer fits the graph adaptively with a small $K$ rather than overfitting it with a large $K$. Squirrel has a large number of heterophilic edges, requiring a more complex frequency response (see Figure LABEL:fig:learned_filters), on which GrokFormer learns to adaptively use a large $K$ for capturing rich frequency components instead of restricting in using small $K$ values. Additionally, we can observe that the maximum order coefficient $\alpha_{k}$ is distributed within the $K<3$ order filter basis on Cora, while on Squirrel, as the order $K$ increases, the largest order coefficient $\alpha_{k}$ extends to the higher-order filter basis. This shows that our method can learn to adaptively assign a large weight to the most suitable filter bases, thereby achieving a synthesized filter specifically for the training graph. More results can be found for the other datasets in Appendix B.3.
@@
Cora Texas PROTEINS IMDB-B
Self-attention-E 77.42$\pm$1.77 90.33$\pm$2.36 71.4$\pm$3.7 74.9$\pm$3.8
Graph Fourier KAN 88.98$\pm$1.25 93.62$\pm$3.04 76.1$\pm$3.5 87.3$\pm$2.5
Full Model 89.57$\pm$1.43 94.59$\pm$2.08 78.2$\pm$2.4 88.5$\pm$2.3
@@ 5.4 Ablation Study
The ablation study is performed to analyze the performance of GrokFormer (Full Model) compared to its two variants: i) Self-attention-E, which contains only efficient self-attention mechanism in GrokFormer, with the proposed spectral graph convolution module removed; ii) Graph Fourier KAN that keeps the spectral graph convolution module only. The results on two node classification datasets and two graph classification datasets are reported in Table 5. It can be observed that Self-attention-E shows good performance in some datasets such as Texas, outperforming most spatial methods in Table 2, due to its ability to capture the feature similarity of the global node. However, it struggles to perform well on the other graph datasets due to its limitation in capturing non-low frequency graph information. The proposed Graph Fourier KAN significantly enhances the performance, showing competitive performance against the state-of-the-art competing methods in Table 2. This superiority benefits from its order and spectrum adaptability that enables expressive graph representation learning without using self-attention. The full model of GrokFormer effectively synthesizes these two modules, achieving consistently better performance than its two variants.
@@ A.1 DATASETS
Datasets Cora Citeseer Pubmed Photo Physics Penn94 Chameleon Squirrel Actor Texas
#Nodes 2,708 3,327 19,717 7,650 34,493 41,554 2,277 5,201 7,600 183
#Edges 5,429 4,732 44,338 238,163 247,962 1,362,229 36,101 217,073 33,544 295
#Features 1,433 3,703 500 745 500 4,814 2,325 2,089 931 1,703
#Classes 7 6 3 8 5 2 5 5 5 5
$\mathcal{H}$ 0.81 0.74 0.80 0.83 0.91 0.47 0.23 0.22 0.22 0.06
@@ A.2.2 Node classification
  • Number of Fourier series expansion terms: {16, 32};
@@ A.2.3 Graph classification
We use the Adam optimizer to train all models, Following (Sun et al., 2020), we perform 10-fold cross validation. We report the average and standard deviation of validation accuracies across the 10 folds within the cross-validation. We implement readout operations by conducting max pooling to obtain a global embedding for each graph. Hyperparameter selection range is as follows:
@@
  • Epoch: {30, 50, 100, 150};
  • Learning rates: {0.001, 0.005, 0.0005, 0.0001};
@@
  • Dropout rate: {0.0, 0.1, 0.2};
  • Batch size: {32, 64, 128};
  • Number of Fourier series expansion terms: {16, 32};
  • Hidden dim: {32, 64};
@@ B.2 VISUALIZATION OF LEARNED FILTERS ON REAL-WORLD DATASETS
In this section, we visualize more filter results learned by our GrokFormer on real-world datasets. From Figure LABEL:learned_filter, we find that (1) on homophilic graphs, our proposed filter learns low-pass filters with different amplitude and frequency responses for them, which is consistent with the homophily property, i.e., the low-frequency information is important in the homophilic scenario. (2) Edges in Chameleon and Penn94 datasets are dense and mostly heterophilic, so our proposed filter learns comb-alike filters with complex frequency components for them. (3) our GrokFormer filter learns an all-pass filter on the Actor dataset protecting its raw features, which is consistent with the fact that its raw features are associated with labels (He et al., 2021). (4) In heterophilic graph Texas, our proposed filter learns a high-pass filter, which matches the strong heterophily, i.e., the high-frequency information that captures the difference between nodes is important.
B.3 ADAPTIVITY IN GRAPH ORDER
@@
Figure LABEL:fig:appendix_order_adaptive shows additional results on order adaptivity. We observe that GrokFormer achieves the best performance when $K$ is small on all homophilic datasets. This is because the low-pass filters desired by homophilic networks (see Figure LABEL:learned_filter) are easy to learn. GrokFormer fits the graph adaptively with a small $K$ rather than overfitting it with a large $K$. In addition, as shown in Figure LABEL:learned_filter, a high-pass filter required by the strong heterophilic network Texas, and an all-pass filter desired by the Actor are also easy to learn, so GrokFormer fits them adaptively with a small $K$. However, for Penn94 and Chameleon, which have dense edges and predominantly heterophilic, complex comb-like filters (see Figure LABEL:learned_filter) are required. As a result, GrokFormer learns to adaptively use a larger $K$ to capture a broader range of frequency components, rather than restricting itself to a small $K$. Moreover, we can find that GrokFormer filter assign the largest order coefficient $a_{k}$ to $K\leq 3$ order filter basis on these datasets that expect simple filters (low-pass, all-pass, high-pass), while on Penn94 and Chameleon, the largest order coefficient $a_{k}$ extends to the higher-order filter basis. This demonstrates that our method can effectively learn order-adaptive filters for datasets with varying properties.
B.4 EMPIRICAL TIME AND SPACE COMPLEXITIES
Cora Penn94
Memory (MB) Times (s) Memory (MB) Times (s)
PolyFormer 1836 13.58 14113 121.78
Specformer 1509 4.35 5053 9.39
GrokFormer 1267 3.23 4647 8.13
Setup. To perform a fair comparison, we run each model for 1,000 epochs and report the total time and space costs. We set the hidden dimension $d$ = 64 for all methods. For Specformer and our GrokFormer, we use use full eigenvectors for Cora and 6,000 eigenvectors for Penn94. We set $K=10$ for polynomial bases of PolyFormer and the order for our GrokFormer. The pre-processing of all models is not included in the training time. The results are shown in Table 8.

### v3

@@ 5 Experiments
In this section, we conduct comprehensive experiments on both synthetic and real-world datasets to verify the effectiveness of our GrokFormer. More experiments can be seen in Appendix B.
@@ 5.1 Performance for Node Classification
Dataset Description. We conduct node classification experiments on 11 widely used datasets in previous graph spectral models (Bo et al., 2023; He et al., 2021; Deng et al., 2024), including six homophilic datasets, i.e., Cora, Citeseer, Pubmed, the Amazon co-purchase graph Photo (He et al., 2021), an extracted subset of Wikipedia’s Computer Science articles–WikiCS (Dwivedi et al., 2023), and a co-authorship network Physics (Shchur et al., 2018; Chen et al., 2024a). We also evaluate on five heterophilic datasets, i.e., Wikipedia graphs Chameleon and Squirrel, the Actor co-occurrence graph (Pei et al., 2020), webpage graphs Texas from WebKB, and Penn94, a large-scale friendship network from the Facebook 100 (Lim et al., 2021). A more detailed description can be found in Appendix A.1.
Baselines and Settings. We compare GrokFormer with sixteen competitive baselines, including four spatial-based GNNs, five spectral-based GNNs, and seven GTs. Note that PolyFormer has multiple variants, and we use the PolyFormer(Cheb) as the baseline. Following the previous works (He et al., 2021; Huang et al., 2024; Bo et al., 2023), we randomly split the node set into train/validation/test set with ratio 60%/20%/20%, and generate 10 random splits to evaluate all models on the same splits. We report the average classification accuracy and standard deviation for each model. For polynomial GNNs, we set the order of polynomials $K$ = 10, consistent with their original setting. For the baseline models, we adopt the hyperparameters provided by the authors. In the large-scale datasets Physics and Penn94, we implement truncated spectral decomposition for both GrokFormer and Specformer to enhance scalability, selecting the 3,000 eigenvectors associated with the smallest (low-frequency) and largest (high-frequency) eigenvalues. More detailed settings can be found in Appendix A.2.
@@
Results. The results are reported in Table 2, which shows the superiority of GrokFormer over state-of-the-art baselines in both homophilic and heterophilic datasets.
Both spatial-based and spectral-based models perform well on the homophilic networks as they can easily capture the similarity information between neighbors, and the low-pass filter is easy to fit in the homophilic networks. Although some GT-based methods, like GraphGPS, also capture the similarity information between neighbors by integrating GNNs with Transformers, resulting in certain performance improvement compared to vanilla Transformer, their heavy emphasis on global information and dependence on a freely learned attention matrix often make them susceptible to overfitting.
@@ 5.2 Performance for Graph Classification
Dataset. We also conduct graph classification experiments on five TU benchmarks from diverse domains. They include three bioinformatics graph datasets, i.e., PROTEINS (Borgwardt et al., 2005), PTC-MR (Toivonen et al., 2003), and MUTAG (Debnath et al., 1991) and two social network datasets, i.e., IMDB-BINARY and IMDB-MULTI (Yanardag & Vishwanathan, 2015) (see Appendix A.1 for more details).
@@ 5.3.1 Pre-defined Filters in Synthetic Datasets
Following prior work (Bo et al., 2023), we generate datasets with six filter patterns of various levels of difficulty. Specifically, images with a resolution of 100 $\times$ 100 from the Image Processing in Matlab library [^1] https://ww2.mathworks.cn/products/image.html are taken, with the image represented as a 2D regular grid graph with 4-neighborhood connectivity. The pixel values serve as node signals ranging from 0 to 1. These image share the same adjacency matrix $\mathbf{A}\in\mathbb{R}^{10000\times 10000}$ and the $m$-th image has its graph signal $x_{m}\in\mathbb{R}^{10000}$. We apply six different predefined filters to the spectral domain of its signal, with each filter detailed in Table 4, where $h_{\delta}(\lambda)$ of Low-comb is defined as $I_{[0,0.5]}(\lambda)+\left|sin(\pi\lambda)\right|I_{(0.5,1)}+\left|sin(2\pi\lambda)\right|I_{[1,2]}$, with $I_{\Omega}=1$ when $\lambda\in\Omega$, $I_{\Omega}=0$ otherwise.
We compare the capability of our GrokFormer filter with six baselines, including GCN, GAT, ChebyNet, GPRGNN, BernNet, and Specformer, in fitting these pre-defined filter patterns through a node regression task. The hyperparameters for our model were probed in $M\in\{16,32,64,128,256\}$, $K\in\{1,2,\cdots,10\}$. For baselines, we set the hyper-parameters suggested by their authors and tune the hidden dimensions to maintain a consistent parameter scale. Two popular evaluation criteria – the sum of squared errors and the $R^{2}$ score – are used.
@@ 5.3.2 Unknown Filters in Real-world Homophilic and Heterophilic Datasets
Adaptability in Graph Spectrum. To analyze the importance of spectrum adaptability, we plot the filters learned by GrokFormer on two typical homophilic datasets, Cora and Citeseer, and two heterophilic datasets, Squirrel and Texas, in Figure LABEL:fig:learned_filters. We observe that GrokFormer learns a low-pass filter on Cora and Citeseer, aligning with their strong homophily. In contrast, on Texas, which exhibits strong heterophily, GrokFormer adaptively learns a high-pass filter to capture the differential information between nodes. On Squirrel, which has dense and mostly heterophilic edges, GrokFormer also effectively learns the heterophily, resulting in a comb-alike filter. Importantly, GrokFormer does not require prior knowledge to manually tune the spectrum hyperparameters to achieve this; it learns to adaptively fit different filters hidden in the graphs with different homophilic and heterophilic properties. Similar results can be found for the other datasets in Appendix B.2.
Adaptability in Graph Spectral Order. Similarly, we also analyze the order adaptability of GrokFormer on Cora, Pubmed, Squirrel and Chameleon. To this end, we evaluate the accuracy performance of GrokFormer with all other settings fixed except that we increase the order $K$ from one to six, from which we observe the relation of $k$ and $K$ w.r.t. the best accuracy. The results are shown in Figure LABEL:fig:k. Cora and Pubmed are strong homophilic network that expect the low-pass filter, and such filter is easy to learn, on which GrokFormer fits the graph adaptively with a small $K$ (i.e., $K\leq 3$) rather than overfitting it with a large $K$. Squirrel and Chameleon have a large number of heterophilic edges, requiring a more complex frequency response (see Figure LABEL:fig:learned_filters), on which GrokFormer learns to adaptively use a large $K$ (i.e., $K>3$) for capturing rich frequency components instead of restricting in using small $K$ values. Additionally, we can observe that the maximum order coefficient $\alpha_{k}$ is distributed within the $K\leq 3$ order filter basis on Cora and Pubmed, while on Squirrel and Chameleon, as the order $K$ increases, the largest order coefficient $\alpha_{k}$ extends to the higher-order filter basis. This shows that our method can learn to adaptively assign a large weight to the most suitable filter bases, thereby achieving a synthesized filter specifically for the training graph. More results can be found for the other datasets in Appendix B.3.
@@
Cora Penn94 Texas PROTEINS IMDB-B
SE 77.42$\pm$1.77 76.29$\pm$0.35 90.33$\pm$2.36 71.4$\pm$3.7 74.9$\pm$3.8
GFKAN 88.98$\pm$1.25 81.36$\pm$0.29 93.62$\pm$3.04 76.1$\pm$3.5 87.3$\pm$2.5
Full Model 89.57$\pm$1.43 83.59$\pm$0.26 94.59$\pm$2.08 78.2$\pm$4.6 88.5$\pm$5.8
@@ 5.4 Ablation Study
The ablation study is performed to analyze the performance of GrokFormer (Full Model) compared to its two variants: i) Self-attention-E (SE), which contains only efficient self-attention mechanism in GrokFormer, with the proposed spectral graph convolution module removed; ii) Graph Fourier KAN (GFKAN) that keeps the spectral graph convolution module only. The results on three node classification datasets and two graph classification datasets are reported in Table 5. It can be observed that Self-attention-E shows good performance in some datasets such as Texas, outperforming most spatial methods in Table 2, due to its ability to capture the feature similarity of the global node. However, it struggles to perform well on the other graph datasets due to its limitation in capturing non-low frequency graph information. The proposed Graph Fourier KAN significantly enhances the performance, showing competitive performance against the state-of-the-art competing methods in Table 2. This superiority benefits from its order and spectrum adaptability that enables expressive graph representation learning without using self-attention. However, on the large-scale dataset penn94, which exhibits a low level of homophily, relying solely on spectral information from Graph Fourier KAN proves insufficient. GrokFormer achieves consistently improved performance only when both Graph Fourier KAN and self-attention are integrated, demonstrating its ability to effectively synthesize the strengths of both modules and outperform its individual variants.
5.5 Empirical Time and Space Complexities
Setup. To perform a fair comparison, we run each model for 1,000 epochs and report the total time and space costs. We set the hidden dimension $d$ = 64 for all methods. For Specformer and our GrokFormer, we use use full eigenvectors for Cora and 6,000 eigenvectors for Penn94. We set $K=10$ for polynomial bases of PolyFormer. The pre-processing of all models is not included in the training time. The results are shown in Table 6.
Dataset Method Memory (MB) Time (s)
Cora PolyFormer 1836 13.58
Specformer 1509 4.35
GrokFormer 1267 3.23
Penn94 PolyFormer 14113 121.78
Specformer 5053 9.39
GrokFormer 4647 8.13
@@ 6 Conclusion and Future Work
Acknowledgments
This research is supported in part by A*STAR under its MTC YIRG Grant (No. M24N8c0103), the Ministry of Education of Singapore under its Tier-1 Academic Research Fund (No. 24-SIS-SMU-008), and the Lee Kong Chian Fellowship.
Impact Statement
This paper presents work whose goal is to advance the field of graph machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 DATASETS
Datasets Cora Citeseer Pubmed Photo WikiCS Physics Penn94 Chameleon Squirrel Actor Texas
#Nodes 2,708 3,327 19,717 7,650 11701 34,493 41,554 2,277 5,201 7,600 183
#Edges 5,429 4,732 44,338 238,163 216,123 247,962 1,362,229 36,101 217,073 33,544 295
#Features 1,433 3,703 500 745 300 500 4,814 2,325 2,089 931 1,703
#Classes 7 6 3 8 10 5 2 5 5 5 5
$\mathcal{H}$ 0.81 0.74 0.80 0.83 0.57 0.91 0.47 0.23 0.22 0.22 0.06
@@ A.2.2 Node classification
  • Number of Fourier series expansion terms: {16, 32, 64};
@@ A.2.3 Graph classification
We use the Adam optimizer to train all models. Following (Sun et al., 2020), we perform 10-fold cross validation. We report the average and standard deviation of validation accuracies across the 10 folds within the cross-validation. We implement readout operations by conducting max pooling to obtain a global embedding for each graph. Since the computational complexity of vanilla self-attention in graph classification task can be alleviated by tuning batch size, we employ the vanilla self-attention mechanism in the graph-level representation learning. Hyperparameter selection range is as follows:
@@
  • Epoch: {100, 200, 300};
  • Learning rates: {0.01, 0.005, 0.001};
@@
  • Dropout rate: {0.0, 0.05, 0.1};
  • Number of Fourier series expansion terms: {16, 32, 64};
  • Hidden dim: {32, 64, 128};
@@ B.2 VISUALIZATION OF LEARNED FILTERS ON REAL-WORLD DATASETS
In this section, we visualize more filter results learned by our GrokFormer on real-world datasets. From Figure LABEL:learned_filter, we find that (1) on homophilic graphs, our proposed filter learns low-pass filters with different amplitude and frequency responses for them, which is consistent with the homophily property, i.e., the low-frequency information is important in the homophilic scenario. (2) Edges in Chameleon and Penn94 datasets are dense and mostly heterophilic, so our proposed filter learns comb-alike filters with complex frequency components for them. (3) our GrokFormer filter learns an all-pass filter on the Actor dataset protecting its raw features, which is consistent with the fact that its raw features are associated with labels (He et al., 2021).
B.3 ADAPTIVITY IN GRAPH SPECTRAL ORDER
@@
Figure LABEL:fig:appendix_order_adaptive shows additional results on order adaptivity. We observe that GrokFormer achieves the best performance when $K$ is small on all homophilic datasets. This is because the low-pass filters desired by homophilic networks (see Figure LABEL:learned_filter) are easy to learn. GrokFormer fits the graph adaptively with a small $K$ rather than overfitting it with a large $K$. In addition, as shown in Figure LABEL:fig:learned_filters and Figure LABEL:learned_filter, a high-pass filter required by the strong heterophilic network Texas, and an all-pass filter desired by the Actor are also easy to learn, so GrokFormer fits them adaptively with a small $K$. However, for Penn94, which have dense edges and predominantly heterophilic, complex comb-like filters (see Figure LABEL:learned_filter) are required. As a result, GrokFormer learns to adaptively use a larger $K$ to capture a broader range of frequency components, rather than restricting itself to a small $K$. Moreover, we can find that GrokFormer filter assign the largest order coefficient $a_{k}$ to $K\leq 3$ order filter basis on these datasets that expect simple filters (low-pass, all-pass, high-pass), while on Penn94, the largest order coefficient $a_{k}$ extends to the higher-order filter basis. This demonstrates that our method can effectively learn order-adaptive filters for datasets with varying properties.
B.4 GRAPH CLASSIFICATION AND REGRESSION
Table 9: Detailed information of additional graph-level datasets.
#Graphs Avg. #nodes Avg. #edges Node feat. (dim) Edge feat. (dim) Tasks Metric
ZINC 12,000 23.2 24.9 Atom Type (28) Bond Type (4) Regression MAE
CIFAR10 60,000 117.6 941.1 Pixel[RGB]+Coord (5) Node Dist (1) Classification ACC
Peptides-func 15,535 150.9 307.3 Atom Encoder (9) Bond Encoder (3) Classification AP
Table 10: Results on additional graph-level datasets. ‘*’ means edge feature is not encoded.
[…]
Graphormer 0.122 - - -
GraphGPS 0.070 72.31 0.6535 0.6257
GRIT 0.060 75.67 0.6988 0.6458
GrokFormer 0.076 74.26 0.6415 0.5987
In this section, we conducts experiments on additional graph-level datasets, including a subset (12K) of ZINC molecular graphs (250K) dataset (Irwin et al., 2012), the super-pixels dataset CIFAR10 (Dwivedi et al., 2023), and a long-range graph benchmark Peptides-func (Dwivedi et al., 2022). We choose graph Transformers with positional or structural embedding (Graphormer, GraphGPS and GRIT) as the baselines. All experiments are conducted on the standard train/validation/test splits of the evaluated benchmarks. We use the hyperparameter for the baselines as suggested in their respective papers. Our hyper-parameters and ranges were as follows:
  • Dropout: {0.0, 0.05, 0.1};
  • Number of layers: {4, 6, 8};
  • Number of Fourier series expansion terms: {16, 32, 64};
  • Number of heads: {1, 2, 3, 4, 5};
  • Learning rate: {0.001, 0.0001, 0.0005};
  • Number of K: {1, 2, 3, 4, 5, 6};
  • Weight decays: {5e-4, 5e-5};
  • Internal MPGNN: {GCN, GatedGCN(Bresson & Laurent, 2017)};
From the results, we observe that GrokFormer achieves better performance than Graphormer and performs comparably well to GraphGPS. GrokFormer slightly underperforms GRIT on ZINC and CIFAR10, while the performance margin on Peptides-func is relatively large. GRIT is designed to improve GT’s expressiveness in large datasets by incorporating graph inductive biases, so it shows an advantage on long-range graph dataset. Although GraphGPS achieves better performance on ZINC and Peptides-func, it exhibits suboptimal results on CIFAR10 and the node-level datasets due to overfitting. In GrokFomer, we aim to enhance the GT’s capability to capture various frequency information on graphs by designing an expressive filter. It achieves a relatively better balance between the generalization ability and the expressiveness, leading to a robust performance on both node-level and graph-level datasets.